OpenStudy (aonz):

Find values of x for (logx)(logx^2) + logx^3 - 5 = 0 Halp please(medal for best answer)

OpenStudy (tyteen4a03):

\((logx)(logx^2) + logx^3 - 5 = 0\) Remember that log(x^y) = y*log(x), then we have (logx)(2logx) + 3logx - 5 = 0 Now you should be able to reduce this equation into a quadratic one.

OpenStudy (aonz):

sorry but i never done surds in brackets which is multiplication. well i dont think i have...

OpenStudy (aonz):

@tyteen4a03 help please

OpenStudy (tyteen4a03):

Surds in brackets? How did you get there?

OpenStudy (aonz):

(logx)(2logx) + 3logx - 5 = 0 These brackets?

OpenStudy (aonz):

i mean logs in brackets :P

OpenStudy (aonz):

@tyteen4a03 HELP PLEASE :(

OpenStudy (aonz):

i completely have no idea what to do... Help please

OpenStudy (tyteen4a03):

logx * logx = log^2(x).

OpenStudy (aonz):



OpenStudy (aonz):